# Activation Approximations Can Incur Safety Vulnerabilities Even in Aligned LLMs: Comprehensive Analysis and Defense

### v1

@@ Abstract
Large Language Models (LLMs) have showcased remarkable capabilities across various domains. Accompanying the evolving capabilities and expanding deployment scenarios of LLMs, their deployment challenges escalate due to their sheer scale and the advanced yet complex activation designs prevalent in notable model series, such as Llama, Gemma, Mistral. These challenges have become particularly pronounced in resource-constrained deployment scenarios, where mitigating inference efficiency bottlenecks is imperative. Among various recent efforts, activation approximation has emerged as a promising avenue for pursuing inference efficiency, sometimes considered indispensable in applications such as private inference. Despite achieving substantial speedups with minimal impact on utility, even appearing sound and practical for real-world deployment, the safety implications of activation approximations remain unclear.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities and gained surging popularity ever since the release of ChatGPT [1]. Their widespread applications across diverse fields are further bolstered by open-source LLMs such as Llama [2], Gemma [3], Mistral [4], Falcon [5], and Qwen [6], offered by model developers capable of conducting extensive pretraining and meticulous alignment procedures like Meta, Google, and Alibaba. Application builders, ranging from enterprises to individuals, can harness the capabilities and flexibility of these open-source aligned LLMs for custom use cases and diverse deployment scenarios. Still, the prohibitive deployment costs hinder the full application potential of LLMs due to their astronomical scales, particularly in resource-constrained scenarios. Additionally, the computational overhead introduced by advanced and complex activation functions like GELU [7] and SwiGLU [8], which are integral and pervasive to modern LLMs (e.g., all above-mentioned open-source LLMs), further exacerbates the inference efficiency challenge. Consequently, improving the inference efficiency of LLMs has garnered growing research attention, with existing approaches exploring distinct facets, including weight quantization [9, 10, 11], weight sparsification [12, 13, 14, 15], model pruning [12, 16, 13], speculative decoding [17, 18]. These efforts offer complementary reductions in inference overhead and hold promise for achieving aggregated efficiency gains when combined.
Activation Approximation. Recently, activation approximation has emerged as a promising and, in some cases, indispensable avenue for pursuing inference efficiency enhancement in LLMs. This line of methods focuses on devising various approximations to alleviate the computation of complex non-linear activation functions, thereby catering to specific deployment requirements. Based on the application scenario and approximation techniques, there are three notable categories, including (i) Activation Polynomialization [19, 20, 21, 22, 23]: Replacing complex nonlinearities in activation functions with polynomials substantially accelerates computation, making it indispensable for the private inference scenario. (ii) Activation Sparsification [24, 25, 26]: Truncating input activation values close to zero so that weight channels corresponding to zero-valued activations can be ignored in inference computation. (iii) Activation Quantization [27, 28]: Quantizing the activation vector from higher bit-widths to lower bit-widths. Table 1 summarizes exemplary inference efficiency gains achieved by representative methods from each category, demonstrating significant wall-clock speed-ups ranging up to 24.6$\times$.
@@
LLM Safety. On par with utility, safety has become an equally paramount dimension for LLMs, garnering widespread consensus from academia, industry, and beyond. Maintaining safety capabilities is critical for preventing LLMs from being exploited to produce harmful responses, such as facilitating unethical uses, spreading misinformation, or fostering discriminatory behaviour. Moreover, maliciously crafted prompts through jailbreak attacks can further induce LLMs towards harmful generations [30, 31, 32], providing an inkling of how intricate the task of safety preservation for LLMs is. In response, model developers have implemented extensive safety alignment procedures, employing techniques such as instruction tuning, reinforcement learning with human feedback (RLHF) [33, 34], and direct preference optimization (DPO) [35], before open-sourcing the aligned LLMs. Unfortunately, safety has turned out to be fragile to maintain, as even minor modifications to the weights of aligned LLMs through fine-tuning can degrade their safety capabilities. When it comes to inference efficiency improvement techniques, safety studies thus far have evaluated weight quantization, weight sparsification, model pruning, and decoding algorithms, demonstrating degraded safety capabilities [36, 37, 38, 39, 40, 41, 42].
@@
Safety Enhancement: Based on these key observations, we propose activation approximation-robust safety alignment, a unified approach to address safety vulnerabilities arising from different activation approximation techniques. Our method enables LLM developers to integrate this additional robustness into their existing safety alignment procedures with minimal modifications and computational overhead (e.g., adding just 2 to 3 lines of code). The resulting LLMs, after being open-sourced and deployed by downstream application providers, remain resilient to various activation approximations introduced to enhance inference efficiency. Experimental results demonstrate that our approach not only handles activation approximations across different methods and various magnitudes but also resists a variety of adaptive jailbreak attacks, such as GCG [46] and AutoDAN [47].
@@
  • In-Depth Safety Analysis: Through analysis across 10 safety-aligned LLMs, we uncover that activation approximations (i) compromise safety before utility, (ii) are most harmful in early layers, and (iii) shift harmful activations into benign spaces to evade detection.
  • Novel Safety Enhancements: We propose an Activation Approximation-Aware Alignment (QuadA) to address these vulnerabilities.
@@ 2.1 Activation Approximations
Activation Polynomialization. In private LLM inference [19, 20, 21, 22, 23], the polynomialization of the activation function is widely used to improve the efficiency of private inference. Private inference is a two-party cryptographic protocol based on secure multi-party computation (MPC) and fully homomorphic encryption (FHE). It can achieve that the server learns nothing about the clients’ input and clients learn nothing about the server’s model except the inference result. In contrast to regular inference in plaintext, operations like Gaussian Error Linear Unit (GELU) and LayerNorm require intensive computation and communication in ciphertext, which usually dominates the total inference latency [48, 49, 50, 51]. To address this problem, the server usually replaces the non-linear functions to MPC/FHE-friendly approximations. MPCFormer[49] replaces GELU($x$) to $0.50.125x^{2}+0.25x^{2}+$ 0.5, BOLT [21] replaces GELU($x$) to a 3-segment polynomial function, BumbleBee[22] and NEXUS [23] replace GELU($x$) to a 4-segment polynomial function.
Activation Sparsification. Recent work has observed that activations in the MLP blocks of LLMs are sparse [52, 26]. This implies that only a few rows (or columns) of the corresponding weight matrices are required for the forward pass. Activation sparsification [24, 25, 26] is an alternative method leveraging sparsity in hidden states. Since the weight channels corresponding to zero-valued activations are not used in computation, speed-up can be realized by selectively omitting these weights during memory transfer. Unlike model pruning techniques, which either remove zero-valued weights from the model or set some weights to zero, activation sparsity exploits the zeros that occur dynamically during runtime, depending on the input.
TEAL [24] and CATS [53] have realized training-free activation sparsity based on magnitude pruning. By pruning low-magnitude, non-salient activations, they realize wall-clock speed-ups of up to 1.9$\times$ at 50% sparsity.
Activation Quantization. Quantization has proven to be a promising technique for mitigating both computational and memory overhead in LLMs. Many post-training quantization methods [9, 10, 54] reduce memory consumption through weight-only quantization, where weights are quantized while activations remain in full precision. To further reduce the computational overhead, recently, many works [27, 28, 55] employs weight-activation quantization which quantizes both weight and activation into low-bit values for the execution of low-bit matrix multiplication. For instance, INT8 quantization of weights and activations can halve the GPU memory usage and nearly double the throughput of matrix multiplications compared to FP16.
To exclude the effect of training and weight quantization on the safety evaluation, we only study post-training quantization (PTQ) and keep weight at full precision. We choose SmoothQuant [27] and OmniQuant [28] as our case studies.
2.2 LLM Safety Evaluations and Alignment
@@
where $\pi_{\theta}$ is the model, $x=(x_{1},x_{2},...,x_{n}),(x_{i}\in\mathcal{V})$ is the context including the prompt, and $y=(y_{1},y_{2},...,y_{n}),(y_{i}\in\mathcal{V})$ is the predicted sequence.
Despite the effectiveness of this mechanism to generate coherent and contextually relevant text, it offers limited control over aligning outputs with human preferences. Consequently, prior work [56, 46, 57, 58, 59] has shown that these models may still produce harmful, biased, or otherwise undesirable content. Particularly in high-stakes domains such as healthcare or finance [60, 61]. Therefore, ensuring safe and aligned outputs is crucial, and this has led to the development of alignment techniques that aim to bridge the gap between a model’s generative objective and the need for responsible outputs [62, 63].
Safety Evaluations. Effective safety evaluation is crucial to understanding and ensuring that LLMs generate outputs that align with human values and ethical guidelines. Existing researches on LLM safety evaluation typically involves multiple dimensions, including Attack Success Rate (ASR) [46], Refusal mechanisms [64], and Safety Risk Index (SRI) [65]. These evaluation metrics have form the backbone of safety assessments in LLM research [66, 67, 68, 69, 70, 71, 72, 73, 74], offering a systematic approach to measuring a model’s robustness against harmful content generation. Despite the thoroughness of these approaches, there remains a gap when it comes to evaluating the impact of activation approximation on safety. Our work seeks to fill this gap by investigating how activation approximation could impact the model’s ability to uphold safety and suggesting strategies to mitigate these associated risks [75, 76, 77, 78].
@@
3 Safety Assessment and Analysis
In this section, we aim to provide a systematic safety assessment and analysis of existing activation approximation methods. We will first introduce the threat model of our study in Section 3.1 and then describe the setup of safety evaluations in Section 3.2. In Section 3.3, we elaborate on the specific activation approximation methods under evaluation and summarize their activation errors, covering seven representative techniques from each of the three categories. Finally, in Section 3.4, we identify three key observations from the safety evaluation results, revealing the characteristics of activation approximations on LLM safety behaviours.
@@ 3.1 Threat Model
Figure 1: Illustration of the three parties in Threat Model.
In this paper, we consider a common scenario in the development and deployment of LLM-based applications, as outlined in Figure 1. This scenario involves three parties: the LLM model developer, the LLM-based application provider, and the application users. In particular, the roles of the first two parties can sometimes overlap within the same organization.
  • LLM Model Developers provide open-source safety-aligned LLMs, enabling application providers to customize models for legitimate usage in accordance with licensing terms. Prominent examples include large technology companies such as Meta (with LLaMA-Chat), Google (with Gemma), and Alibaba (with Qwen), as well as open-source organizations/groups that have released safety-aligned LLMs like Falcon, Mistral, Alpaca, and Zephyr.
@@
  • LLM-based Application Providers customize open-source safety-aligned LLMs to develop applications tailored to specific use cases, expecting the customized models to retain both the utility and safety capabilities of the original LLMs. These providers aim to deliver high-quality, responsive services to their users while striving to minimize inference latency. However, they could face resource constraints in terms of computation and memory. Consequently, they are incentivized to adopt inference efficiency-enhancing techniques to reduce resource usage and wall-clock inference time while maintaining utility and safety capabilities as much as possible.
@@
  • LLM-based Application Users interact with LLM-based applications expecting helpful, accurate, and timely responses. Some might occasionally, for example out of curiosity, issue queries that should be turned down by a safety-aligned LLM. Meanwhile, there could be adversarial users launching jailbreak attacks [46, 47, 89] to deliberately bypass the safety guardrails of the application providers’ LLMs.
@@
Adaptive Capabilities. This paper considers an adaptive adversary who can construct the jailbreak prompt based on gradients of the target LLM [46, 47, 89]. These methods pad a prefix or suffix to the original prompt, which can be optimized to achieve the attack objective. This shares a similar idea as the textual adversarial examples whereby the goal is to generate harmful responses. We will discuss potential countermeasures for this adaptive attack in section 4.2.
3.2 Safety Assessment Setup
@@
Safety Measurement. We measure LLM’s safety by evaluating the Attack Success Rate (ASR), which is defined as the percentage of failure to abstain from responding to a malicious instruction. These malicious instructions come from AdvBench [46], which contains 520 items in total. Previous works [46, 80] simply use a small set of rejection phrases such as “Sorry, I can’t”, “I can not fulfill” to do prefix matching to determine whether the model rejects the answer. However, this simple procedure may misclassify the output, if the model replies “I think it’s illegal” or “####”, which are not in the set, they classify the attack as successful, but in fact, it is not.
@@
However, we found that it could still fail to distinguish a benign output from a misaligned one, especially under circumstances where the instructions are complicated. We show an example of a false positive produced by the HarmBench classifier. This example was obtained from DRA (USENIX Security’24) [90] where they proposed a jailbreak attack that bypasses safety checks by hiding a harmful question in a word puzzle, and asking the LLM to solve the puzzle and then answer the question. Their prompt-based jailbreak approach gives relatively complicated instructions for the LLMs to extract the questions and answer in a step-by-step manner. In the case of the example response, it is clear that the LLM did not produce a concrete and actionable plan to fulfill the attacker’s malicious request. However, due to the "positive" answer, the HarmBench classifier became confused and incorrectly identified the LLM’s response as harmful.
To ensure the validity of a defense mechanism, it is crucial to have a misalignment evaluation metric that can accurately reflect the effectiveness of the proposed approach. Based on the observed existence of false positives, we cannot directly rely on HarmCLS to produce a precise account of the resulting ASR of the LLMs after implementing our defense strategy. We collect 120 additional prompt-response pairs from the DRA [90] and manually mark whether it is harmful or not. We fine-tuned the HarmBench classifier [64] with this dataset to achieve a higher accuracy for misbehaviour judgment.
Utility Measurement. We use perplexity on WikiText-2 [91] to measure the model’s utility, which is one of the most common utility metrics. Perplexity is defined as the exponentiated average negative log-likelihood of a sequence. For a tokenized sequence $X=(x_{1},...,x_{n})$, the perplexity of $X$ is:
@@
(a) Error distributions in activation polynomialization, with BOLT [21] as the example.
(b) Error distributions in activation sparsification, with TEAL [24] in 50% sparsity.
(c) Error distributions in activation quantization, with SmoothQuant [27] in 8-bit activation.
Figure 2: Activation error distributions on Llama-3.1-8B-Instruct using different activation approximations. The first column presents errors before $\mathbf{W}^{\text{up}}$, and the second column presents errors before $\mathbf{W}^{\text{down}}$. Best-fit Gaussian/Laplace distributions are overlaid in black for comparison.

### v2

@@ Abstract
Large Language Models (LLMs) have showcased remarkable capabilities across various domains. Accompanying the evolving capabilities and expanding deployment scenarios of LLMs, their deployment challenges escalate due to their sheer scale and the advanced yet complex activation designs prevalent in notable model series, such as Llama, Gemma, Mistral. These challenges have become particularly pronounced in resource-constrained deployment scenarios, where mitigating inference bottlenecks is imperative. Among various recent efforts, activation approximation has emerged as a promising avenue for pursuing inference efficiency, sometimes considered indispensable in applications such as private inference. Despite achieving substantial speedups with minimal impact on utility, even appearing sound and practical for real-world deployment, the safety implications of activation approximations remain unclear.
@@
Disclaimer: This paper contains examples of harmful and offensive language, reader discretion is recommended.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities and gained surging popularity ever since the release of ChatGPT [1]. Their widespread applications across diverse fields are further bolstered by open-source LLMs such as Llama [2], Gemma [3], Mistral [4], Falcon [5], and Qwen [6], offered by model developers capable of conducting extensive pretraining and meticulous alignment procedures like Meta, Google, and Alibaba. Application builders, ranging from enterprises to individuals, can harness the capabilities and flexibility of these open-source aligned LLMs for custom use cases and diverse deployment scenarios. Still, the prohibitive deployment costs hinder the full application potential of LLMs due to their astronomical scales, particularly in resource-constrained scenarios. Additionally, the computational overhead introduced by advanced and complex activation functions like GELU [7] and SwiGLU [8], which are integral and pervasive to modern LLMs (e.g., all above-mentioned open-source LLMs), further exacerbates the inference efficiency challenge. Consequently, improving the inference efficiency of LLMs has garnered growing research attention, with existing approaches exploring distinct facets, including weight quantization [9, 10, 11], weight pruning [12, 13, 14, 15], speculative decoding [16, 17]. These efforts offer complementary reductions in inference overhead.
Activation Approximation. Recently, activation approximation has emerged as a promising and, in some cases, indispensable avenue for pursuing inference efficiency enhancement in LLMs. This line of methods focuses on devising various approximations to alleviate the computation of complex non-linear activation functions, thereby catering to specific deployment requirements. Based on the application scenario and the approximation techniques, there are three notable categories, including (i) Activation Polynomialization [18, 19, 20, 21, 22]: Replacing complex nonlinearities in activation functions with polynomials substantially accelerates computation, making it indispensable for the private inference scenario. (ii) Activation Sparsification [23, 24, 25]: Truncating input activation values close to zero so that weight channels corresponding to zero-valued activations can be ignored in inference computation. (iii) Activation Quantization [26, 27]: Quantizing the activation vector from higher bit-widths to lower bit-widths. Table 1 summarizes exemplary inference efficiency gains achieved by representative methods, demonstrating significant wall-clock speed-ups with activation approximations.
@@
LLM Safety. On par with utility, safety has become an equally paramount dimension for LLMs, garnering widespread consensus from academia, industry, and beyond. Maintaining safety capabilities is critical for preventing LLMs from being exploited to produce harmful responses, such as facilitating unethical uses, spreading misinformation, or fostering discriminatory behaviour. Moreover, maliciously crafted prompts through jailbreak attacks can further induce LLMs towards harmful generations [29, 30, 31], providing an inkling of how intricate the task of safety preservation for LLMs is. In response, model developers have implemented extensive safety alignment procedures, employing techniques such as instruction tuning, reinforcement learning with human feedback (RLHF) [32, 33], and direct preference optimization (DPO) [34], before open-sourcing the aligned LLMs. Unfortunately, safety has turned out to be fragile to maintain, as even minor modifications to the weights of aligned LLMs through fine-tuning can degrade their safety capabilities [35, 36].
LLM Safety Studies on Inference Efficiency Improvement Techniques. In addition to the activation approximations mentioned above, various other lines of techniques have also explored different components of the LLM inference pipeline to improve inference efficiency, such as weight quantization [37, 9] and pruning [12, 38], context pruning [39] and compression [40], KV Cache management [41, 42] and KV compression [43, 44], speculative decoding [16, 45], and so on. Among these, existing safety investigations have primarily focused on weight quantization and pruning techniques [46, 47, 48]. Hong et al.[46] assess the safety performance of compressed LLMs, with a focus on weight quantization and pruning, and demonstrate that weight approximation techniques degrade the safety capabilities of aligned LLMs. Similarly, Belkhiter et al.[47] report consistent findings, observing safety degradation in weight-quantized LLMs. Egashira et al.[48] reveal that the weight quantization process can be exploited to produce a harmful quantized LLM. While these studies focus exclusively on the safety concerns of weight approximation, the safety issues of various activation approximation techniques remain largely unexplored, including activation polynomialization, activation sparsification, and activation quantization. Notably, although the activation quantization methods to be investigated in this paper also employ quantization, they are applied to activations rather than to weights.
@@
Safety Enhancement: Based on these key observations, we propose activation approximation-aware safety alignment, a unified approach to address safety vulnerabilities arising from different activation approximation techniques. Our method enables LLM developers to integrate this additional robustness into their existing safety alignment procedures with minimal modifications. The resulting LLMs, after being deployed by downstream application providers, remain resilient to various activation approximations introduced to enhance inference efficiency. Experimental results demonstrate that our approach not only handles activation approximations across different methods and various magnitudes but also resists a variety of adaptive jailbreak attacks.
@@
  • In-Depth Safety Analysis: Through analysis across 10 safety-aligned LLMs, we uncover that activation approximations (i) compromise safety before utility, (ii) are most harmful in early few layers, and (iii) shift harmful activations into benign spaces to evade detection.
  • Novel Safety Enhancements: We propose an Activation Approximation-Aware Alignment (QuadA) to address these vulnerabilities. Our code has been open-sourced at https://github.com/Kevin-Zh-CS/QuadA.
@@ 2 Background
Figure 1: Illustration of the three parties in the threat model: the LLM developer, the LLM-based service provider, and the user.
@@ 2.1 Activation Approximations
Activation Polynomialization. In private LLM inference [18, 20, 21, 22], the polynomialization of the activation function is widely used to improve the efficiency of private inference. Private inference is a two-party cryptographic protocol based on secure multi-party computation (MPC) and fully homomorphic encryption (FHE). It can achieve that the server learns nothing about the clients’ input and clients learn nothing about the server’s model except the inference result. In contrast to regular inference in plaintext, operations like Gaussian Error Linear Unit (GELU) and LayerNorm require intensive computation and communication in ciphertext, which usually dominates the total inference latency [52, 53]. To address this problem, the server usually replaces the non-linear functions to MPC/FHE-friendly approximations. MPCFormer[52] replaces GELU($x$) to $0.50.125x^{2}+0.25x^{2}+$ 0.5, BOLT [20] replaces GELU($x$) to a 3-segment polynomial function, BumbleBee[21] and NEXUS [22] replace GELU($x$) to a 4-segment polynomial function.
Activation Sparsification. Recent work has observed that activations in the MLP blocks of LLMs are sparse [54, 25]. This implies that only a few rows (or columns) of the corresponding weight matrices are required for the forward pass. Activation sparsification [23, 25] is an alternative method leveraging sparsity in hidden states. Since the weight channels corresponding to zero-valued activations are not used in computation, speed-up can be realized by selectively omitting these weights during memory transfer. Unlike model pruning techniques, which either remove zero-valued weights from the model or set some weights to zero, activation sparsity exploits the zeros that occur dynamically during runtime, depending on the input. TEAL [23] and CATS [55] have realized training-free activation sparsity based on magnitude pruning. By pruning low-magnitude, non-salient activations, they realize wall-clock speed-ups of up to 1.9$\times$ at 50% sparsity.
Activation Quantization. Quantization has proven to be a promising technique for mitigating both computational and memory overhead in LLMs. Many post-training quantization methods [9, 10] reduce memory consumption through weight-only quantization, where weights are quantized while activations remain in full precision. To further reduce the computational overhead, recently, many works [26, 27] employs weight-activation quantization which quantizes both weight and activation into low-bit values for the execution of low-bit matrix multiplication. For instance, INT8 quantization of weights and activations can halve the GPU memory usage and nearly double the throughput of matrix multiplications compared to FP16. To exclude the effect of training and weight quantization on the safety evaluation, we only study post-training quantization (PTQ) and keep weight at full precision. We choose SmoothQuant [26] and OmniQuant [27] as our case studies.
2.2 LLM Safety Alignment
@@
where $\pi_{\theta}$ is the model, $x$ is the context including the prompt, and $y=(y_{1},y_{2},...,y_{n}),(y_{i}\in\mathcal{V})$ is the predicted sequence.
@@
3 Assessment and Analysis
@@ 3.1 Threat Model
In this paper, we consider a common scenario in the development and deployment of LLM-based applications, as outlined in Figure 1. This scenario involves three parties: the LLM developers, the LLM-based service providers, and the service users. Based on the threat model, this paper reveals that activation approximations employed by LLM-based service providers to improve inference efficiency can undermine the safety capabilities of LLMs, even when the models have been safety-aligned by LLM developers, thereby incurring severe safety implications for a broad range of service users.
  • LLM Developers perform extensive safety alignment during post-training to provide safety-aligned LLMs, enabling LLM-based service providers to customize the models for legitimate use in line with licensing terms. Typically, LLM Developers include Meta, Mistral AI, DeepSeek, and Alibaba. They utilize safety datasets during post-training to align their models, aiming to prevent them from being induced to generate harmful content via jailbreak attacks.
@@
  • LLM-based Service Providers customize safety-aligned LLMs to develop applications tailored to specific use cases, expecting the customized models to retain both the utility and safety capabilities of the original LLMs. These providers aim to deliver high-quality, responsive services while striving to minimize inference latency. However, they could face resource constraints in terms of computation and memory. Consequently, they are incentivized to adopt inference efficiency-enhancing techniques to reduce resource usage and wall-clock inference time while maintaining utility and safety capabilities as much as possible.
@@
  • Users interact with LLM-based applications expecting helpful, accurate, and timely responses. Some might occasionally, for example out of curiosity, issue queries that should be turned down by a safety-aligned LLM. Meanwhile, there could be adversarial users launching jailbreak attacks [60, 75, 76] to deliberately bypass the safety guardrails of the application providers’ LLMs.
@@
Adaptive Capabilities. In addition to jailbreaking with pure harmful queries, this paper also considers an adaptive adversary who can construct the jailbreak prompt based on gradients of the target LLM [60, 75, 76]. These methods pad a prefix or suffix to the original prompt, which can be optimized to achieve the attack objective. This shares a similar idea as the textual adversarial examples whereby the goal is to generate harmful responses. We will discuss potential countermeasures for this adaptive attack in Section 5.1.
3.2 Assessment Setup
@@
Safety Measurement. We measure LLM’s safety by evaluating the Attack Success Rate (ASR) and the Harmful Score (HS). ASR is defined as the percentage of failure to abstain from responding to a malicious instruction. These malicious instructions come from AdvBench [60], which contains 520 items in total. Previous works [60, 57] simply use a small set of rejection phrases such as “Sorry, I can’t”, “I can not fulfill” to do prefix matching to determine whether the model rejects the answer. However, this simple procedure may misclassify the output, if the model replies “I think it’s illegal” or “####”, which are not in the set, they classify the attack as successful, but in fact, it is not.
@@
However, we found that it could still fail to distinguish a benign output from a misaligned one, especially under circumstances where the instructions are complicated. We show an example of a false positive produced by the HarmBench classifier in Appendix A.4. This example was obtained from DRA (USENIX Security’24) [77] where they proposed a jailbreak attack that bypasses safety checks by hiding a harmful question in a word puzzle, and asking the LLM to solve the puzzle and then answer the question. Their prompt-based jailbreak approach gives relatively complicated instructions for the LLMs to extract the questions and answers in a step-by-step manner. In the case of the example response, it is clear that the LLM did not produce a concrete and actionable plan to fulfill the attacker’s malicious request. However, due to the "positive" answer, the HarmBench classifier became confused and incorrectly identified the LLM’s response as harmful. To ensure the validity of a defense mechanism, it is crucial to have a misalignment evaluation metric that can accurately reflect the effectiveness of the proposed approach. Based on the observed existence of false positives, we cannot directly rely on HarmCLS to produce a precise account of the resulting ASR of the LLMs after implementing our defense strategy. We collect 120 additional prompt-response pairs from the DRA [77] and manually mark whether it is harmful or not. We fine-tuned the HarmBench classifier [74] with this dataset to achieve a higher accuracy for misbehaviour judgment.
Table 2: Safety and utility assessment metrics
Metric Dataset Size Range
Safety ASR $\downarrow$ AdvBench 520 [0,100]
HS $\downarrow$ BeaverTails 8,210 [1,5]
Utility PPL $\downarrow$ WikiText-2 245,569 >1
MMLU $\uparrow$ MMLU 15,908 [0,100]
For Harmful Score (HS), we follow established practice [35] to utilize GPT-4 to judge the harmfulness automatically. We conduct the evaluations on the BeaverTails [78] dataset, which contains 8,210 harmful questions spanning 14 categories. Specifically, we construct a prompt comprising the prohibited model usage policy, the harmful input instruction, the model’s output, and a rating rubric. This prompt is then submitted to GPT-4 via the OpenAI API, instructing the model to assess whether the output violates the specified usage policy. Further details of the scoring methodology can be found in Appendix A.3.
Utility Measurement. We utilize WikiText-2 [79] and the Massive Multitask Language Understanding (MMLU) benchmark [80] to measure the model’s utility. For the WikiText-2 benchmark, we employ perplexity as the utility metric, which is defined as the exponentiated average negative log-likelihood of a sequence. For a tokenized sequence $X=(x_{1},...,x_{n})$, the perplexity of $X$ is:
@@
Massive Multi-task Language Understanding (MMLU) [80] is a popular benchmark for evaluating the utility of LLMs. MMLU consists of 15,908 multiple-choice questions, spanning across 57 subjects from highly complex STEM fields and international law to nutrition and religion. It is one of the most widely used benchmarks and is prevalently leveraged by mainstream literature [2, 6, 56, 81]. For the MMLU benchmark, we employ the overall average accuracy as the utility metric. We summarize the safety and utility assessment metrics in Table 2.